Gabriella and Steve have adjusted gross incomes of $47,100 and $32,700, respectively. Assume that each person takes one exemption and the standard deduction. Answer parts ​(a) through ​(c) below

Tax Rate Single Married Filing Jointly
10% up to $8,925 up to $17,850
15% up to $36,250 up to $72,500
25% up to $87,850 up to $146,400
28% up to $183,250 up to $223,050
Standard Deduction $6,100 $12,200
Exemptions​ (per person) $3,900 $3,900

a. Calculate the tax owed by the couple if they delay their marriage until next year so they can each file a tax return at the single tax rate this year.

The couple owes $___ (Simplify your answer. Round to the nearest dollar as​ needed.)

b. Calculate the tax owed by the couple if they marry before the end of the year and file a joint return.

The couple owes $___

c. Does the couple face a​ "marriage penalty" if they marry before the end of the​ year?

yes or no
Answer #1

Total Income, of both the person = 47,100 + 32, 700 = 79,800

A. If they both files their return as an individual, then their individual taxes are 5,789 and 3,154 respectively. Hence as a couple they both owes 8,943.

B. If they file a joint return, then their tax would be 8,468.

C. No they won't be facing any "Marriage Penalty" if they marry before the end of the year. Because, if they marry before the end of the year, then thy would be paying less tax in contrast of what they would pay as an individual. The difference amount is (8,943 - 8,468 = 475).

NOTE: Since, the tax calculation are not asked in the question, hence we are not showing them here. However, the above values have been calculated after making the calculations.